OpenStudy (ilovepurple42043):

what is the value of tan (-2pi/3)

iYuko (iyuko):

Like 1.7

OpenStudy (er.mohd.amir):

tan(-2pi/3)=-tan(2pi/3)=-tan 120=-tan(180-60)=-*-tan 60=tan60 =sqrt3=1.732

OpenStudy (welshfella):